Two alternative start-up FinTech projects were being contemplated for financing by a venture capitalist to determine which one is more viable, based on cost and returns. Table 4.1 below shows a five-year schedule for the two projects:

Table 4.1
Project Start of Project End of 1st Year End of 2nd Year End of 3rd Year End of 4th Year End of 5th Year
BlockChain GoPay ($230,000) ($32,000) $66,000 $96,000 $106,000 $119,000
DLT CloudPay ($276,000) $20,000 $65,000 $96,000 $102,000 $118,000

(a) If you were the Project Manager on the Venture Capitalist team, using the Net Present Value (NPV) method, which project would you recommend be financed based strictly on the schedule shown above and an interest rate of 7.5%?
(b) Use a Microsoft Excel, or any other method to deduce/calculate the Internal Rate of Return (IRR) for both projects. Explain how you would advise the Bank which project to finance using the result from the IRR method?
1
Expert's answer
2020-04-29T09:28:03-0400

(a) NPV for BlockChain GoPay is:

"NPV = \sum{CF/(1 + r)^n} = -230,000/1.075^0 - 32,000/1.075^1 + ... + 119,000/1.075^5 =550,891.63."

NPV for DLT CloudPay is:

"NPV = \sum{CF/(1 + r)^n} = -276,000/1.075^0 - 20,000/1.075^1 + ... + 118,000/1.075^5 =34,699.06."

Hence, the project Block chain GoPay should be accepted since the NPV is higher than that of DLT cloudPay.

(b) IRR is an interest rate at which NPV = 0.

Using function "IRR" in Excel we find that IRR for project BlockChain GoPay IRR is 12% and for project GoPay is 11%.